OpenStudy (anonymous):

S and T are related by the following relationship . S=k/T , where k is a constant. If S increases by 50% by how much does T decreases? and vice versa if T increases by 50% how much does S decreases?

OpenStudy (anonymous):

it is suppose to be 33.33% but I don;t know why

OpenStudy (anonymous):

S=k/T means ST=k is also true, and k is constant. Suppose S increases by 50% then the new is is 3/2 times as large as the old S. In order for the equation (3/2)ST=k to be true, the new T must be (2/3) the old T, which means T is decreasing by (1/3)T, or 33.33333%.

OpenStudy (anonymous):

Similarly If T decreases by 50%, then the new T is 1/2 times the old T, so S(1/2)T = k needs to have a new S of 2S in order to be true. In order to become 2S, S must double, which is the same as increasing by 100%.

OpenStudy (anonymous):

@cebroski ah yes I get it now. thank you !